Oil flows through a standard 25 mm diameter orifice under 5.5 m head at a


measured rate of 3 L/s. The jet strikes a wall 1.5 m away and 0.12 m


vertically below the centreline of the contracted section of the jet. Determine


all the coefficients of the orifice.

1
Expert's answer
2022-03-21T17:57:25-0400

"C_v=\sqrt{\frac{x^2}{4yH}}=0.92,"

"C_d=\frac Q{A\sqrt{2gH}}=0.59,"

"C_c=\frac{C_d}{C_v}=0.64."